40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 36 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  851   Tue Feb 28 22:50:11 2012 FrankNotesScheduleupdate

The following tasks will be finished within the next 24h:

  • aligning both cavities
  • align beat
  • install RF-AM PD (start with one to get estimate for noise budget)
  • measure TF from seismic to beat (if possible, vacuum chamber temp might not be stable enough)

we will then focus on (re-)measuring some things we need very soon for a more accurate NB (there are more but the following ones are my favorites for now. We are about a factor of 3 above the CTN @200Hz)

  • RF-AM - not in NB yet, but seems to dominate 100Hz-1kHz region, so we do this first (seen last week while optimizing everything)
  • RIN  - old measurement must be wrong as it is too high
  • seismic  - new stack, so new TF
  • phase noise of Marconi locked to rubidium clock - we know that the noise will be less below 1kHz, above it most likely does not change the phase noise level
  • electronic noise from both PDH loops


  850   Tue Feb 28 00:19:41 2012 FrankNotesDAQdaqd restarted

restarted the framebuilder process. Channel list needs to be updated, i didn't clean up the FSS channels so there are still old/unused channel names.

Current list of saved channels.

106 channels total
105 trend channels

chnum   slow    |name                           |rate   |trend  |group  |bps    |bytes  |offset |type   |active
0       0       C3:PSL-BOX_SENS1                                16      1       0       4       64      0       4       1
0       0       C3:PSL-BOX_SENS2                                16      1       0       4       64      64      4       1
0       0       C3:PSL-BOX_SENS3                                16      1       0       4       64      128     4       1
0       0       C3:PSL-BOX_SENS4                                16      1       0       4       64      192     4       1
0       0       C3:PSL-BOX_TEMPAVG                              16      1       0       4       64      256     4       1
0       0       C3:PSL-BOX_HEATER                               16      1       0       4       64      320     4       1
0       0       C3:PSL-BOX_SETPT                                16      1       0       4       64      384     4       1
0       0       C3:PSL-BOX_KP                                   16      1       0       4       64      448     4       1
0       0       C3:PSL-BOX_KI                                   16      1       0       4       64      512     4       1
0       0       C3:PSL-BOX_KD                                   16      1       0       4       64      576     4       1
0       0       C3:PSL-BOX_ENABLE                               16      1       0       4       64      640     4       1
0       0       C3:PSL-BOX_TIMEOUT                              16      1       0       4       64      704     4       1
0       0       C3:PSL-BOX_SCALE                                16      1       0       4       64      768     4       1
0       0       C3:PSL-ACAV_RFPDDC                              16      1       0       4       64      832     4       1
0       0       C3:PSL-ACAV_RCTRANSPD                           16      1       0       4       64      896     4       1
0       0       C3:PSL-ACAV_LOCKEDLEVEL                         16      1       0       4       64      960     4       1
0       0       C3:PSL-RCAV_SENS1                               16      1       0       4       64      1024    4       1
0       0       C3:PSL-RCAV_SENS2                               16      1       0       4       64      1088    4       1
0       0       C3:PSL-RCAV_SENS3                               16      1       0       4       64      1152    4       1
0       0       C3:PSL-RCAV_SENS4                               16      1       0       4       64      1216    4       1
0       0       C3:PSL-RCAV_TEMP                                16      1       0       4       64      1280    4       1
0       0       C3:PSL-RCAV_TEMPAVG                             16      1       0       4       64      1344    4       1
0       0       C3:PSL-RCAV_S1CAL                               16      1       0       4       64      1408    4       1
0       0       C3:PSL-RCAV_S2CAL                               16      1       0       4       64      1472    4       1
0       0       C3:PSL-RCAV_S3CAL                               16      1       0       4       64      1536    4       1
0       0       C3:PSL-RCAV_S4CAL                               16      1       0       4       64      1600    4       1
0       0       C3:PSL-RCAV_SETPT                               16      1       0       4       64      1664    4       1
0       0       C3:PSL-RCAV_KP                                  16      1       0       4       64      1728    4       1
0       0       C3:PSL-RCAV_KI                                  16      1       0       4       64      1792    4       1
0       0       C3:PSL-RCAV_KD                                  16      1       0       4       64      1856    4       1
0       0       C3:PSL-RCAV_ENABLE                              16      1       0       4       64      1920    4       1
0       0       C3:PSL-RCAV_TIMEOUT                             16      1       0       4       64      1984    4       1
0       0       C3:PSL-RCAV_SCALE                               16      1       0       4       64      2048    4       1
0       0       C3:PSL-RCAV_RFPDDC                              16      1       0       4       64      2112    4       1
0       0       C3:PSL-RCAV_RCTRANSPD                           16      1       0       4       64      2176    4       1
0       0       C3:PSL-RCAV_LOCKEDLEVEL                         16      1       0       4       64      2240    4       1
0       0       C3:PSL-FSS_HEATER                               16      1       0       4       64      2304    4       1
0       0       C3:PSL-FSS_FREQCOUNT                            16      1       0       4       64      2368    4       1
0       0       C3:PSL-FSS_VCOMON                               16      1       0       4       64      2432    4       1
0       0       C3:PSL-FSS_VCOMON_CAL                           16      1       0       4       64      2496    4       1
0       0       C3:PSL-FSS_VCOFREQ                              16      1       0       4       64      2560    4       1
0       0       C3:PSL-FSS_RFPDDC                               16      1       0       4       64      2624    4       1
0       0       C3:PSL-FSS_LODET                                16      1       0       4       64      2688    4       1
0       0       C3:PSL-FSS_PCDET                                16      1       0       4       64      2752    4       1
0       0       C3:PSL-FSS_FAST                                 16      1       0       4       64      2816    4       1
0       0       C3:PSL-FSS_PCDRIVE                              16      1       0       4       64      2880    4       1
0       0       C3:PSL-FSS_RCTLL                                16      1       0       4       64      2944    4       1
0       0       C3:PSL-FSS_VCODET                               16      1       0       4       64      3008    4       1
0       0       C3:PSL-FSS_TIDALOUT                             16      1       0       4       64      3072    4       1
0       0       C3:PSL-FSS_MODET                                16      1       0       4       64      3136    4       1
0       0       C3:PSL-FSS_VCODETPWR                            16      1       0       4       64      3200    4       1
0       0       C3:PSL-FSS_MIXERM                               16      1       0       4       64      3264    4       1
0       0       C3:PSL-FSS_SLOWM                                16      1       0       4       64      3328    4       1
0       0       C3:PSL-FSS_VCOM                                 16      1       0       4       64      3392    4       1
0       0       C3:PSL-FSS_TIDALINPUT                           16      1       0       4       64      3456    4       1
0       0       C3:PSL-FSS_SW1                                  16      1       0       4       64      3520    4       1
0       0       C3:PSL-FSS_SW2                                  16      1       0       4       64      3584    4       1
0       0       C3:PSL-FSS_PHFLIP                               16      1       0       4       64      3648    4       1
0       0       C3:PSL-FSS_VCOTESTSW                            16      1       0       4       64      3712    4       1
0       0       C3:PSL-FSS_VCOWIDESW                            16      1       0       4       64      3776    4       1
0       0       C3:PSL-FSS_INOFFSET                             16      1       0       4       64      3840    4       1
0       0       C3:PSL-FSS_MGAIN                                16      1       0       4       64      3904    4       1
0       0       C3:PSL-FSS_FASTGAIN                             16      1       0       4       64      3968    4       1
0       0       C3:PSL-FSS_PHCON                                16      1       0       4       64      4032    4       1
0       0       C3:PSL-FSS_RFADJ                                16      1       0       4       64      4096    4       1
0       0       C3:PSL-FSS_SLOWDC                               16      1       0       4       64      4160    4       1
0       0       C3:PSL-FSS_VCOPWR                               16      1       0       4       64      4224    4       1
0       0       C3:PSL-FSS_VCOMODLEVEL                          16      1       0       4       64      4288    4       1
0       0       C3:PSL-FSS_TIDALSET                             16      1       0       4       64      4352    4       1
0       0       C3:PSL-FSS_LOCK                                 16      1       0       4       64      4416    4       1
0       0       C3:PSL-FSS_SLOWLOOP                             16      1       0       4       64      4480    4       1
0       0       C3:PSL-PMC_PMCTLL                               16      1       0       4       64      4544    4       1
0       0       C3:PSL-PMC_RFPDDC                               16      1       0       4       64      4608    4       1
0       0       C3:PSL-PMC_LODET                                16      1       0       4       64      4672    4       1
0       0       C3:PSL-PMC_PMCTRANSPD                           16      1       0       4       64      4736    4       1
0       0       C3:PSL-PMC_PCDRIVE                              16      1       0       4       64      4800    4       1
0       0       C3:PSL-PMC_PZT                                  16      1       0       4       64      4864    4       1
0       0       C3:PSL-PMC_MODET                                16      1       0       4       64      4928    4       1
0       0       C3:PSL-PMC_PMCERR                               16      1       0       4       64      4992    4       1
0       0       C3:PSL-PMC_SW1                                  16      1       0       4       64      5056    4       1
0       0       C3:PSL-PMC_SW2                                  16      1       0       4       64      5120    4       1
0       0       C3:PSL-PMC_PHFLIP                               16      1       0       4       64      5184    4       1
0       0       C3:PSL-PMC_BLANK                                16      1       0       4       64      5248    4       1
0       0       C3:PSL-PMC_GAIN                                 16      1       0       4       64      5312    4       1
0       0       C3:PSL-PMC_INOFFSET                             16      1       0       4       64      5376    4       1
0       0       C3:PSL-PMC_PHCON                                16      1       0       4       64      5440    4       1
0       0       C3:PSL-PMC_RFADJ                                16      1       0       4       64      5504    4       1
0       0       C3:PSL-PMC_RAMP                                 16      1       0       4       64      5568    4       1
0       0       C3:PSL-PMC_LOCK                                 16      1       0       4       64      5632    4       1
0       0       C3:PSL-PEM_RMTEMP                               16      1       0       4       64      5696    4       1
0       0       C3:PSL-PEM_BOXTEMP                              16      1       0       4       64      5760    4       1
0       0       C3:PSL-FSS_RFAM_RCAV                            16      1       0       4       64      5824    4       1
0       0       C3:PSL-FSS_RFAM_ACAV                            16      1       0       4       64      5888    4       1
0       0       C3:PSL-FSS_EOM_TSET                             16      1       0       4       64      5952    4       1
0       0       C3:PSL-FSS_EOM_TACT                             16      1       0       4       64      6016    4       1
0       0       C3:PSL-FSS_EOM_IMON                             16      1       0       4       64      6080    4       1
0       0       C3:PSL-FSS_EOM_SETTEMP                          16      1       0       4       64      6144    4       1
0       0       C3:PSL-GEN_DAQ1                                 16      1       0       4       64      6208    4       1
0       0       C3:PSL-GEN_DAQ2                                 16      1       0       4       64      6272    4       1
0       0       C3:PSL-GEN_DAQ3                                 16      1       0       4       64      6336    4       1
0       0       C3:PSL-GEN_DAQ4                                 16      1       0       4       64      6400    4       1
0       0       C3:PSL-GEN_DAQ5                                 16      1       0       4       64      6464    4       1
0       0       C3:PSL-GEN_DAQ6                                 16      1       0       4       64      6528    4       1
0       0       C3:PSL-GEN_DAQ7                                 16      1       0       4       64      6592    4       1
0       0       C3:PSL-GEN_DAQ8                                 16      1       0       4       64      6656    4       1
10001   0       C3:FB1-FB_DUMMY                                 16384   0       0       4       65536   6720    4       0

  849   Mon Feb 27 23:57:25 2012 FrankNotesDAQchannel list for framebuilder updated

changed the channel list for epics channels which will be recorded. Modified file according to changes mentioned in previous post.
File contains also channels which physically don't exist at the moment, e.g. RF-AM channels and 2nd. temp stabilized box around chamber.

File: /cvs/cds/caltech/chans/daq/C3PSL_EPICS.ini

  848   Mon Feb 27 21:45:55 2012 FrankNotesRefCavnew beam height for cavities

The beam height changed by 1/8". Current beam height is 5-7/8". The top stack plate is slightly off-centered towards the ACAV side but we don't want to re-open the chamber now to fix this (it's not much, but one can see it). We can do this the next time we open it anyway or if we see weird coupling between different stack modes. When we open it next time we will add some markings at the end faces of the top stack plate to better see if it's centered or not.

We will align the beams along the hole pattern as designed first and move the cavities to the right position. Once we add the air springs (hopefully soon) we have to change the beam height again.

  847   Mon Feb 27 20:52:42 2012 FrankDailyProgressDAQDAQ reconfigured, channels added/removed

cleaned up the VME stuff a little bit - removed all old channels we don't use anymore, mostly FSS stuff, added new channels for the in-vac sensors and renamed a bunch of others to match the current situation. The only thing i didn't touch is all channels from "RCAV", which is now common to both cavities but i didn't want to dig too deep and change all scripts etc. So that's left for the near future.

Current channels connected to 16bit ADC (VMIC 3123):

  • CH0 : C3:PSL-PEM_BOXTEMP (Sensor in 2nd thermal box around vacuum chamber)
  • CH1 : C3:PSL-PEM_RMTEMP (Sensor measuring temp in clean enclosure)
  • CH2 : C3:PSL-RCAV_TEMP (4 Sensors on vacuum chamber minus voltage reference)
  • CH3 : C3:PSL-RCAV_SHIELD1 (In-Vac sensor on RCAV copper shield)
  • CH4 : C3:PSL-ACAV_SHIELD1 (In-Vac sensor on ACAV copper shield)
  • CH5 : C3:PSL-ACAV_SHIELD2 (In-Vac sensor on ACAV copper shield)
  • CH6 : C3:PSL-FSS_FREQCOUNT (Frequency counter output measuring beat)
  • CH7 : C3:PSL-FSS_VCOMON (VCO feedback monitor signal)
  • CH8 : C3:PSL-BOX_SENS1 (Sensor #1 of 2nd thermal box around vacuum chamber)
  • CH9 : C3:PSL-BOX_SENS2 (Sensor #2 of 2nd thermal box around vacuum chamber)
  • CH10 : C3:PSL-BOX_SENS3 (Sensor #3 of 2nd thermal box around vacuum chamber)
  • CH11 : C3:PSL-BOX_SENS4 (Sensor #4 of 2nd thermal box around vacuum chamber)
  • CH12 : C3:PSL-RCAV_SENS1 (Sensor #1 on vacuum chamber)
  • CH13 : C3:PSL-RCAV_SENS2 (Sensor #2 on vacuum chamber)
  • CH14 : C3:PSL-RCAV_SENS3 (Sensor #3 on vacuum chamber)
  • CH15 : C3:PSL-RCAV_SENS4 (Sensor #4 on vacuum chamber)

new/modified channels are available in real-time, but not saved at the moment!

old database-files are located in "20120227"-subfolder on the SUN in the psl folder and on the svn in /software

  846   Mon Feb 27 19:17:57 2012 taraNotesElectronics EquipmentNoise in delay line read out technique

I estimate the noise sensitivity in the delay line technique. The calculation should tell us how good this method can be used to measure frequency noise in beat signal.

 The calculation follows the setup in PSL:828. All relevant datasheets can be found here.


      I try to compute how each component in the setup generates noise and shows up at the end of the stream. With the calibration factor, I can convert the noise back to its equivalent frequency noise at the input.

      The only component that introduces noise in the setup is the ZHL-1A mixer. Its noise figure is ~ 8dB. Assuming that the only noise from the input side is thermal noise in 50 ohm, then the noise level after the amplifier is ~ 50nV, see details in the note below.

     For each signal trace, the signal goes through 4-ch splitter (-6dB), cable, then the mixer (~3dB conversion loss). These components give a factor of (1/4) x (1/2) to the signal. I'm not sure how noise in both delay line will sum up at the mixer. So for now I just assume noise coupling from one side of the mixer. The noise level after the low pass should be 50nV/8 = 6.3 nV. The calibration from Voltage to frequency noise at the mixer output is 2.5 MHz/V (from measurement on 2012_02_22,160MHz, svn)  Thus the absolute frequency noise  is  6.2nV x 2.5 MHz/V = 15.5 mHz.  This level is ~ a factor of 1.5 lower than the frequency noise of Marconi (10kHz tuning range) , see psl:834, psl:833. If this is the only limiting source, we should be able to measure coating noise upto ~500Hz (limited bandwidth due to the chosen delay time is not taken into account yet).

     It would be nice to be able to measure the noise and compare it with the calculation. To measure the noise, we need a low noise input source (Marconi with 1kHz tuning range should be ok) with power as specified in the setup. However, the measured noise level is higher than the expected noise and we don't know what the cause is. So we can not verify the calculated noise level yet.


Noise figure = 10log10( Noise Factor). From the datasheet, noise figure of the amplifier ~ 8dB which corresponds to Noise Factor = 6.3.

Noise Factor = SNR from input / SNR from output. For our setup, the signal from input is 5dBm (0.4 V), with noise ~ 1nV flat (50 Ohm thermal noise). the output signal is 22.8dBm (3.1V). Thus the expected noise at the output = Noise Factor x (signal_out/signal_in) x noise_in = 6.3x (3.1/0.4) x 1nV ~ 50nV.

  845   Mon Feb 27 10:56:43 2012 taraDailyProgressopticbeam realignment after new seismic stack installation

I realigned ACAV and found TEM00, but now the transmitted beam is completely missed the opening on the insulation, it is off from the center by ~ 1 cm.

  844   Sun Feb 26 17:40:43 2012 taraDailyProgressopticbeam realignment after new seismic stack installation

I'm trying to re-align the beams to the cavities. Due to the new RTV springs for the seismic stack, the cavities' natural axes shift by ~1/4 " with respect to the previous position.

     I had to adjusted the height of the top mirror of the periscope before I could align and lock RCAV (visibility ~ 95%) again. The pictures below show the position of the current beam. With the previous setup, the beam position was almost at the center of the holes. Now, for RCAV, the axis shifts closer to the edge. RCAV might yaw with respect to the previous position. Left picture shows the incoming beam position, Right picture shows the outgoing beam position.


      For ACAV, however, it seems that the position changes a lot and the beam clips on the outer edge of the top mirror before I can even find TEM00. I think I'll have to add a spacer between the mirror mount and the vertical plate in order to re align the beam.

     I think we can keep the stack position as it is for now, if I can lock both cavities and the transmitted beams can be adjusted on the breadboard for beat path. We might also have to increase the hole size on the insulation cap as well depending on where the beam position of ACAV will be.


  843   Sun Feb 26 02:18:38 2012 FrankNotesDAQnew channels for in-vac sensors

the new AD590 sensors are physically connected to

  • VME3123, C0 S3 (RCAV)
  • VME3123, C0 S4 (ACAV #2)
  • VME3123, C0 S5 (ACAV #1)

measured across 30kOhm (+/-1%, 100ppm) to gnd, so voltage is ~9V at 30C. Had no precision resistors but doesn't matter as the channels are used for monitoring purposes only

Software channels are not assigned yet.

  842   Sun Feb 26 01:46:09 2012 FrankDailyProgressVacuumturbo disconnected - ion pump enabled

disconnected the turbo pump and enabled the ion pump  - initial current was 1mA  (value before venting was <0.1uA)

  841   Sat Feb 25 01:32:39 2012 FrankNotesRefCavin-vacuum connector pinout

pinout of 9-pin PEEK in-vac D-SUB connector:

1(-) / 6(+) - AD590 no1 (ACAV)
2(-) / 7(+) - AD590 no2 (ACAV)
3(-) / 8(+) - AD590 no3 (RCAV)
5 / 9 - heater (155Ohms) (ACAV)


  840   Fri Feb 24 21:59:44 2012 Tara, FrankDailyProgressVacuumupgrades installed - pumping started

we've installed:

  • the new stack -see earlier post with measured TF
  • both copper radiation shields
    - one shield has a NiChrome heater wrapped around (the other one not as we first want to see if we need more/less power, but calculation says this one should be right)
    - three temp sensors (AD590), two on the shield with heater, one on the one without
    - those sensors won't be used for temp control, only for temp and gradient monitoring and/or safety shutdown (if required)
  • cable from D-SUB feedthrough to rack breakout terminal

window is back on and pumping down over the weekend. Thermal insulation for vacuum chamber is back in place, so we should be ready to go for a new measurement on Monday afternoon after installation of the beat breadboard

P1820701.JPG P1820756.JPG

  839   Fri Feb 24 14:39:15 2012 FrankNotesNoiseBudgetthings to do on a very short timescale
  • measure Marconi phase noise LOCKED to Rubidium clock for different input ranges and LO frequencies (make a list of possible frequencies)
  • re-measure coupling from RIN into beat
  • re-measure seismic coupling to beat - we have the stack TF and the coupling factor for RCAV so far
  • measure RF-AM coupling - not limiting at the moment but we don't know where it's gona be
  838   Fri Feb 24 13:01:46 2012 taraNotesNoiseBudgetNoise budget for beat measurement by PLL

I added another noise budget,"beat_nb.m", on SVN. This noise budget is for beat measurement via PLL. It calculates both cavities' noise sources which will add up incoherently in the beat measurement. The code is still incomplete. More noise source will be added later. This noise budget is supposed to be the correct nb for our beat measurement.


This code is based on the previous one which is the noise budget of a single cavity.  I remove a few traces e.g. noises from ambient temperature fluctuation coupling through thermal expansion. Since they are way smaller than the other. I still keep Brownian noise in substrates and spacers. These might be removed later so that the plot is not too crowded.  


The current code includes noise from:

  1. Spacer Brownian (x2 from both cavities)
  2. Substrate Brownian (x2 from both cavities, x2 from two substrates on each cavity)
  3. Coating Brownian   ( x2 cavities, x2 substrates)
  4. Substrate Thermoelastic (x2 caviteis, x2 substrates)
  5. shot noise from PDH locking(x2 from acav and rcav loop, they can be independently adjusted)
  6.  LO phase noise from PLL
  7. PLL read out noise

The other noise sources that will be added later are:

  1. Seismic coupled noise
  2. RIN induced noise

      For noise from seismic, especially from vertical direction, we can measure the TF between seismic -> beat. This will tell us the real coupling. Koji and Frank measured the TF between the table and the stack in the previous entry, so there should be no problem with the setup.  We can also try other directions (horizontal), but from vertical direction, we should be able to get the idea how seismic couples into beat signal.

     For RIN induced noise, I'm not quite sure yet if the SNR will be large enough to be able to measure, but we can approximate.


  837   Fri Feb 24 04:10:41 2012 KojiSummarySeismicNew stack transfer function / ringdows

[Frank and Koji]

1) Stack vertical transfer functions

We have attached a KISTLER accelerometer on the stack.
The accelerometer was epoxied on a low-outgassing kapton tape while the tape is attached on the stack.

The table was shook by either a PZT or impact-hammering by a fist.

The new resonant freqs are 10.5 and 35Hz in stead of previous 15Hz and 55Hz.
This provides us an additional isolation by factor of ~10 above 20Hz

2) Stack pitching mode identification

Location of the accelerometer on the stack was swept from the center to the edge of the plate.
The difference of the transfer functions gives us the idea where are thepitch resonances.
It seems that the resonant frequencies in pitch are 20 and 60Hz

3) Ringdown measurement of the other modes

In order to check the resonant freq of the other modes, the stack was excited by a finger
in longitudinal, yaw-rotational, and transverse directions.

The results: longitudinal 3.1Hz, rotational 5.2Hz, transverse 2.7Hz


Attachment 1: vertical_tf_comparison.pdf
Attachment 2: vertical_pitch_tf.pdf
Attachment 3: ringdown_time_series.pdf
  836   Fri Feb 24 00:03:25 2012 Koji, FrankDailyProgressSeismiccharacterized old and new stack

after carefully thinking about options how to get better results and closer to the coating thermal noise within the next 10 days we decided to open the vacuum chamber and work on the seismic isolation. The the current sensitivity limit is kind of flat and seems to continue like that towards lower frequencies (which we can kind of see when floating the table).

As we could not clearly identify other sources which are limiting us at the moment we decided to improve the seismic isolation next, which might help us measuring CTN at lower frequencies where it is higher. In parallel we will also add the thermal shields and the heater. So we can reduce the drift between the cavities which then makes it possible to reduce the range of the Marconi and so lower the contribution of phase noise. We already locked it to the Rubidium clock which also lowers the phase noise and should make it possible to see CTN below 1kHz.

With the heater we can also tune the beat frequency towards lower frequencies which

  1. increases the SNR on the beat PD as we operate it currently beyond the specs (125MHz max)
  2. lowers the phase noise contribution from the PLL LO (gets lower with lower frequencies)

We first replace the old springs of the stack with the new ones characterized here. Measurements will be posted in a separate entry.

In parallel we will work on the air springs to isolate the whole chamber. But we won't get those parts until mid/end next week so it will be kind of a last-minute change before the LSC meeting.

Things finished so far:

  • replaced RTV springs
  • characterized old and new stack
  • mounted sensors to shields
  • added copper tubes to stack

To-do list:

  • make PTFE part to clamp cables to top stack plate
  • finish wire connections to feedthrough and check everything
  • close chamber and pump down
  • re-align cavities
  • modify base plates for beat breadboard
  • replace one mirror mount base on breadboard




  835   Thu Feb 23 13:40:00 2012 Koji, FrankDailyProgressNoiseBudgetperformance of tuned experiment

- incoming -

We did optimize the alignment, power levels etc and tweeked almost every knob of the system to get an idea where we have to look for the current limit in sensitivity. Didn't find anything dominant. A lot of already known things can limit if you intentionally make it worse / misalign things, but once optimized they are below the current measured noise performance.  Detail later.


Tue Feb 28 21:28:53 2012 

Beat measurement after optimization, floated table.


     The beat noise is roughly a factor of 2 above the coating noise at 130 Hz. This gives us a good reason to change the springs for the seismic stack in order to get better sensitivity at lower frequency, as it is getting closer to the coating noise at lower frequency.  At 2kHz and above, the noise spectrum's feature is similar to the noise budget, but with some offset. We might miss a few other flat noise sources( noise from RFPD, electronics) that we have to add into the noise budget. Most of the mechanical peaks around 100 - 1kHz are probably from the mirror mounts.

  834   Tue Feb 21 13:15:23 2012 FrankDailyProgressBEATMarconi noise with delay line but no amplifier (@160MHz and @20MHz)

once again measured the Marconi noise with the delay line - this time without the amplifier (so using a 7dBm mixer instead of 13dBm) and at two different frequencies, 20MHz and 160MHz. Still have no clue where the flat noise floor is coming from which we've seen in previous measurements (see elog #833).

The measurement at 20MHz (left graph) was taken with the frequency tuned so that the DC offset is close to zero (0.1mV). The measurements show a consistent 1/sqrt(f) noise level at low frequencies, independent from the marconi phase noise. And again, the Marconi noise for 1k input range can't be measured.

The right graph shows the result at 160MHz, but this time with a slight DC offset, so that LO AM couples into the measurements. The slope is similar to the one seen at 20MHz, but with clearly more features which come from amplitude noise of the Marconi.  The situation at high frequencies is the same, the phase noise of 10k range can be seen, for lower input ranges not.

noise_budget_20M.png noise_budget_160M.png

Next i measured the coupling from AM into the signal for different DC offsets (only at 160MHz), this time again for the original mixer and power levels (the one we want to use in our setup). As before at zero DC offset we are insensitive to AM and the noise floor is somewhat flat. With increasing offset the coupling from AM into the measurement becomes more dominant and looks identical to the coupling which can be seen on the upper right plot. This looks similar at other frequencies but i didn't save those.

The question now is: If the (almost) white noise floor is not thermal noise, amplifier noise etc. and not from AM, where does it come from? Any ideas?
I've tried the following thing, but nothing worked:

  • using DC-blocks at different locations
  • using isolation transformers at different locations
  • using several combinations of LP and HP filters
  • different power levels for LO and RF
  • lots of combinations of the above


  833   Fri Feb 17 14:54:36 2012 FrankSummaryBEATmarconi noise measured using delay line

tried to measure the frequency noise of the Marconi using the delay line. Setup is identical to the schematic posted in entry #832.

I've set the LO power to 13.64dBm as it is close to optumum value. The mixer output is terminated with 500Ohms. The slope is 1.1145MHz/V.
Measured the noise at the output using the SR785 and a SR560, gain 1000.

Plot shows the following:

  • SR560, gain=1000 with 50Oms input termination
  • SR560, gain=1000 with 1.9MHz LP-filter and 500Ohms load impedance
  • Mixer output with RF input terminated with 50Ohm, LO still ON
  • Marconi noise with modulation input turned off
  • Marconi noise measured using the delay-line technique (with mod input range 10kHz, turned ON but input terminated with 50Ohms) - noisefloor NOT subtracted
  • Marconi noise measured using the PLL (with mod input range 10kHz, turned ON but input terminated with 50Ohms)


  • SR560 is currently not limiting the performance
  • The noise floor even with FM tuning input turned OFF is too high, but far above the rest of the equipment noise floor
  • We can barely see the additional phase noise from the Marconi with 10k input range turned ON - no way to see the noise of input ranges 1k or below at the moment. Reason unknown. Noise level is higher as predicted from the ideal setup (see elog entry #825)


Attachment 2: noise_budget.fig
  832   Thu Feb 16 23:27:50 2012 Frank, TaraNotesBEAToptimized cable-delay setup sensitivity

Started characterizing the cable-delay setup with the right length of cable (134ft of RG58 for 160MHZ). After checking the change in sensitivity with load impedance i've changed the load to 500 Ohms (instead of the usual 50 Ohms). I think an additional low-impedance path for the 2f has to be put in parallel later (to have proper 50Ohms @ 2f) to not get it reflected at the input of the low-pass filter back into the IF port of the mixer. (see first schematic).



However, the following simple setup has been used for the measurements:



I've measured the output signal vs different LO power levels while keeping the RF signal strength constant (8.29dBm) to find out the optimum signal strength in terms of size (not noise at this point!).
The following plots show the result:

  • The peak mixer output signal (for pi phase shift) increases with LO power
  • at the same time the slope around the zero-crossing shows a maximum around the specified LO power value (13dBm) - update: had a typo in vertical units, now correct - plot itself was correct

LO_power.png LO_power2.png


  831   Thu Feb 16 19:44:23 2012 Frank, TaraNotesBEATmixer output scaling with load impedance


==Some intros about this measurement:==

We are using a mixer as a phase detector. Usually, mixers are not optimized to be used as phase detectors (that provide DC output), they do have 50 ohms impedance output. However, when we use a mixer as a phase detector, the output is DC, and the output behaves like a current source. We want to know how load impedance will affect the readout system, so we measured the output and varied the load impedance.


==Setup and Result==

The setup is shown here. We checked how much we can gain in the cable-delay readout scheme by using a different (higher) load impedance. Used the standard setup with a Marconi as the source, modulated the frequency with a triangular signal and looked after the LP filter with the scope for different termination impedances after the mixer. Plot is in arbitrary units as only the relative change is relevant. Signals are triggered at slightly different times.

mixer_slope.png mixer_slope2.png


 The result shows that with higher load impedance, we have better sensitivity (steeper slope). This is as expected from a current source with load impedance ( V = IR). At this point we are using 500 Ohm load impedance in our regular setup because we are not sure if the higher load impedance will introduce any extra noise.

  830   Thu Feb 16 02:40:58 2012 frank, taraDailyProgressBEATBreadboard setup for beat path

We still working on the breadboard setup. There will be several things we have to modify for the setup.

Now we can mount the mirrors on the three screw blocks, with beam on the hole pattern.

[add details about what to modify]





Note: I have been looking at frequency only around 100 - 3 kHz. Here is the beat for broadband with other noise. The seismic noise in the plot is for floated table, I'll edit that, but it is pretty similar for what we had before. So the breadboard setup has not introduced any noise at low frequency that shows up in the beat yet.


  829   Wed Feb 15 14:09:34 2012 FrankDailyProgressBEATreduced RG58 cable length to optimum value

reduced our RG58C/U cable length to optimum value (134.2ft) and characterized it. Below the confirmation that it is what it should be.

loss is 8.6dB, delay 206.4ns


  828   Tue Feb 14 23:35:04 2012 FrankNotesBEATproposed (new) cable delay setup



Mixer will be driven very hard to saturate it. To operate the mixer in the required saturated mode, the RF signal level should be at least:

  • +1 dBm for Standard Level (+7 dBm LO) mixers
  • +7 dBm for Standard Level (+13 dBm LO) mixers
  • +10 dBm for High Level (+17 dBm LO) mixers

so if we use the right (optimum) cable we would have ~8dBm, which should be perfect for a level 13 mixer.

Let's see if we can confirm the calculations...

  827   Tue Feb 14 22:47:45 2012 frank, taraDailyProgressBEATBreadboard setup for beat path

I checked the mechanical peaks in breadboard setup. We get rid of the peak at 800 Hz from the periscope and a big peak around 200 Hz. However, there are some new peaks popping up which are not identified yet.


   As the beams are not on the whole pattern, Frank suggested that I move the whole board and clamp it down instead. In order to do that, I had to remove the QWPs. I clamped it down with steel clamps around the legs. After that I inserted some damping posts (see below pictures). It turned out that the rubber damping does not help much at this point the noise spectrum does not change at all between with or without the rubber. I think it is because I cannot insert enough rubber between the board and the post, as I slid them in after I clamped the board. I should have placed the damping posts in their places then clamped down the board.

[add clamping fig]




fig: beat noise in different setup: Blue: beat path with periscope, Magenta: beat with breadboard setup, Green: with damping on mirror mounts. The data were taken with 2kHz tuning range, gain 200.


Since only the beat path that has been changed, all the peaks that popping out are contribution from the breadboard setup.

So things that are changed are:

  •      Mirror mounts: all mirror mounts on the bread board for beat path (except ones for picking up light for CCD/PD) are all the same, (the mount with whole frame around the mirror). This causes several peaks cluster around 1-1.2 kHz. By placing rubber cones on those mounts and the peaks were damped down by a factor of 3 or so. So we might need to insert some foam for better damping later.IMG_0332.JPG
  •       Heavy lead blocks are placed on the board to  help reducing noise at lower frequency around 100-200 Hz. Better damping/clamping for the board have to be done.IMG_0328.jpg
  •      The broad band noise was gone after all the stray beams were properly blocked with razor blade beam block.

There are a few new peaks due to the breadboard setup which have not been identified yet. It is very hard to check, since tapping with slight force already excite the peaks of the mirror mounts around 1 kHz. Once the mounts are damped, other peaks might be easier to be found.

  826   Tue Feb 14 21:21:19 2012 FrankDailyProgressBEAToptimum cable length for delay line & comparison with LP-filter

now as we know that the optimum loss of the delay line is 8.68dB we can calculate the optimum cable length.


optimum length for 160MHz are:

  • RG58C/U :    40.9m (134.2ft),     delay= 205.74ns      $0.30/ft
  • RG142B/U :  52.88m (173.5ft),    delay= 251.17ns      $2.25/ft
  • RG316 :        24.86m (81.6ft),     delay=120.81ns       $1.05/ft
  • RG405 :        30.68m (100.6ft),   delay=146.97ns       $6.00/ft
  • LMR-240 :     84.94m (278.7ft),   delay=337.20ns       $0.72/ft
  • LMR-400 :   166.56m (546.5ft),   delay=652.92ns       $0.90/ft

cables which introduce more delay for the same amount (8.68dB) of loss are better.

Now, we compare the minicircuits low-pass filter SLP-200 (datasheet) with the cables.

  • Insertion loss @150MHz is 0.4dB
  • group delay ~6ns

so we could add 22 filters for an optimum total delay/loss ratio. Total group delay would be 132ns.
If we compare now with the delays we get from the cables we see that even the simple RG58 gives us 50% more delay for the same loss ( and the price for the cable is the same as a single filter).
Using RG142 instead we get almost a factor of 2 more sensitivity and even more using lower loss cables.

So i don't see an advantage using those LP filters instead of cables.


  825   Tue Feb 14 17:27:58 2012 FrankDailyProgressBEATfirst (basic) calculations for cable delay line

started with some simple calculations for replacing the PLL with a delay line. Started with modeling the loss in the cable depending on frequency and length (separate matlab-function for different cables).
Below some first plots for our current "situation" (which probably changes in the near future but that's what it is right now)  having a beat note @ 160MHz , 5dBm from the PD and an ZHL-1A amplifier (16dB gain) and a 4-way splitter (for two delay lines with different cable length):

  • loss in cable vs length
  • delay vs length
  • detection range (180deg phase shift) vs cable length
  • signal strength at mixer output vs cable length in dBm and Vpp
  • signal size at mixer output for two different power levels going into the cable
     - the first one using the ZHL-1A amplifier (16dB gain)
     - a second situation where we have 27dBm (0.5W) going into the cable (1W max for a single 2-way power splitter)
  • frequency noise sensitivity for the current situation using the ZHL-1A amplifier and assuming a 1nV/rtHz amplifier at the mixer output (no other noise sources so far!) or an SR560 with 4nV/rtHz.

files are on the svn in "CTNLab\simulations\noise_budget\delay_line_readout".

The optimum sensitivity is reached when the decrease in output signal is compensated by the increase in 2*pi*tau, which happens with a total loss of 8.68dB (factor 1/e) of the cable.

We don't win with adding delay if we make the cable longer, even if we increase the power going into the cable! That also explains why i had such a poor sensitivity with the 500ft of RG58 (which had 33dB of loss). Using 15ft of cable instead would have given the same sensitivity!

UPDATE 2/14/2012@8pm- files on the SVN contain now also data for RG405 and LMR-400 !

cable_loss.png cable_delay.png mixer_range.png

mixer_output_dBm.png mixer_output_Vpp.png mixer_output_sensitivity.png

mixer_output_sensitivity2.png final_sensitivity.png  final_sensitivity_SR560.png

  824   Tue Feb 14 02:39:53 2012 frank, taraDailyProgressBEATBreadboard setup for beat path

We removed the periscopes in beat path and use breadboard setup instead. There are higher broadband noise in the beat around 100 - 3kHz. At least, the peak around 800 Hz is slightly small, since the contribution from the periscope in the beat path was removed.


     We are trying to get rid of individual mechanical peaks in the beat signal. One of the major peaks comes from the periscopes and the associated mirror mounts. So instead of using periscopes to bring the beam height down, we use breadboard setup to bring the whole beat path up. This new setup gets rid out the periscopes, and 4 mirror mounts on it.

    Frank has a solid work assembly file showing how the setup should look like. However, the beam coming out of the cavities are not exactly on the whole pattern, and without the periscope we have no way to steer the beam to the designed path. As a temporary solution, I use a post mounts which are clamped on the board, not screwed.


 Fig1: The beam path is not on the whole pattern on the table, so the block mirror mount cannot be used.


fig 2: current breadboard setup. The mirrors behind quarter waveplates are mounted on a regular post and clamped down on the board. If I use the block mount, the beam won't hit the mirror.  Note: I use the previous beam splitter post (with 1.4kHz resonant peak) for now, because we don't have anything that fits right now.




fig3: Comparison between original setup (in blue,with periscope) and breadboard setup (green). Both signals were taken with 2kHz tuning range, gain 200.


I think the broadband noise might come from the scattering on the PD. THe spotsize on the PD is not much smaller than the PD. I haven't found the appropriate lens for modematching yet.  The peaks around 1kHz-2 kHz also seem to be more than the regular setup. This will be investigated.

  823   Mon Feb 13 13:42:43 2012 FrankNotesBEATphase/group delay for RG58 (and other cables)

measured the delay for the old cable (RG58):  dPhi=180deg, df=600KHz 

1.67ns/ft                (value from datasheet:  1.53ns/ft)


typical values for other cables using the following dielectric materials:

Dielectric Type                           Time Delay (ns/ft)
Solid Polyethylene (PE)               1.54
Foam Polyethylene (FE)              1.27
Foam Polystyrene (FS)                1.12
Air Space Polyethylene (ASP)     1.15-1.21
Solid Teflon (ST)                           1.46
Air Space Teflon (AST)                 1.13-1.20

  822   Mon Feb 13 13:16:02 2012 FrankNotesBEATtemp coefficient of coax cables

from   Measurements of Earth-station delay instabilities using a delay-calibration device  measured at 70MHz  


Group delay [ppm/K] Phase delay [ppm/K]
RG58U -80 68
FSJ1-50 7 -7
ISOCORE 21 -14

other data (e.g. for RG223/U) can be found here : http://tmo.jpl.nasa.gov/progress_report/42-99/99E.PDF

or here : http://ivs.nict.go.jp/mirror/meetings/v2c_wm1/phase_stability.pdf

or here: http://tesla.unh.edu/courses/ece758/Handouts/cable-specs.pdf

Type of cable             Temp coef (ppm/K)
RG-8A                          -85
RG-214                        -60
RG-223                        -40 to -100
Belden 9913               -21
Andrew FSJ4-50B      -2 to +6
Andrew LDF2-50        -8 to +6
Andrew LDF4-50A     +7 to +16


And for the LMR-240 which i would buy for future cable delay lines ~24ppm/K



  821   Sun Feb 12 19:40:50 2012 FrankDailyProgressSeismicre-measuring seismic coupling to cavities

trying again to measure the coupling from seismic to cavity length for the individual cavities. Measuring the coupling to differential length (beat signal) is not a problem at all, but to the individual length.

The problem arises from the small coupling coefficient from vertical acceleration to changes in length. to not measure any effect from the filtering from the stack we have to measure at low frequencies. The first mechanical resonances occur at around 6Hz, the stack itself has it's first resonance at 16Hz. So we have to measure below 5Hz. The coupling to changes in length is small, about 1e-9 * Length of the cavity  [units: m/(m/s2)], so about 2e-10 m/(m/s2). The signal from shaking the table around 1Hz is estimated to be a few hundreds of Hz/rtHz with maximum modulation. However the laser frequency noise is 10kHz/rtHz @1Hz, so we have to integrate very long to get a reasonable SNR. We cant use anything on the table to pre-stabilize the laser to reduce it's noise as this would be shaked as well and we don't really know which one we actually measure.

For the first cavity we have to look at the feedback to the laser pzt as this tells us how much the laser frequency has to be corrected. We assume that shaking the table at 1Hz does not mechanically modulate the laser frequency in any other way. For the second cavity we can't simply lock it the usual way as we would have two coupled cavities (that's what we measure using the beat signal already). So we have to lock the laser to the second cavity instead without using the FSS path by feeding back to the laser (fast actuator) instead of the VCO.

A first measurement showed that we have an additional mechanical resonance around 6Hz which we currently don't have in our current stack model (and actually don't know exactly where it's coming from).
We measured the Eigenmodes of the stack some time ago and have two candidates for it (see here)

  1. beam line, translational motion, f = 6.96 Hz, Q = 21.5 
  2. horizontal transverse motion, f = 6.35Hz, Q = 25.9.

So i will re-measure the TF below 10Hz to clearly identify which one it is.


resonance frequencies of the stack : 16.1Hz and 55.6Hz
other resonance frequency: 6.5Hz

coupling to cavity length: COMSOL model: 53 kHz / (m/s2)
~100 kHz / (m/s2)

All data and plots on the svn in /measurements/2012_02_12.


Notes for calibration:

TF to beat signal:

  • accelerometer sensitivity: 1023mV/g
  • range of PLL VCO: 10kHz -> 7kHz/V

TF to fast actuator:

  • accelerometer sensitivity: 1023mV/g
  • FAST MON output:  3.07 MHz/V


  820   Sun Feb 12 12:21:09 2012 FrankDailyProgressBEATpeak hunting

Did a little bit of peak hunting to clear our frequency span of interest from those massive mechanical resonances we currently have. After replacing the combining beam splitter mount we got rid of the 1.4kHz peak already. Yesterday i've focused on the mounts within the beat setup, but not the periscope, as we already know that this is very unstable and we will take care of that soon. I didn't want to replace things, just know where which stuff comes from.

I've found (only) one mirror mount which is currently clearly visible in our noise spectrum . Tapping the other mounts or damping the front plate or springs does not change the spectrum (at least i don't see any changes). Tapping (even slightly) is very difficult anyway as you also excite all the mounts surrounding your DUT, especially the periscopes and your whole spectrum changes and it's hard to figure out which is your primary resonance you are looking for. So i prefer damping it with a large piece of rubber and than compare it with a spectrum taken before with a reasonable.amount of averages.

Anyway, i found only one mirror mount (out of six) which i could clearly identify in our current noise spectrum. It's one of the mirrors right in front of the combining beam splitter.
Below a comparison before and after damping the front plate of the mirror mount. Resonance frequency is 544Hz. I have to check but i think we can replace this one with a non-adjustable turning mirror.

We still don't know where the 1.1KHz stuff is coming from.


  819   Fri Feb 10 23:47:22 2012 Tara, FrankDailyProgressBEATbeam splitter mount replaced

we replaced the mount for the combining beam splitter in the beat setup as it caused a large, broadband peak in the spectrum around 1.4kHz. The new mount is one of the old, fixed turning mirror blocks they used in initial LIGO at LLO as far as i know. After replacing the mount the peak is entirely gone. I've used two springs instead of one to increase the pressure.  We could not determine the resonance frequency of the new mount. Tapping the mount excites only known mechanical resonances from the surrounding mirror mounts. Tara posted a plot for comparison before and after replacing that mount (see here). He also has prepared a nice plot combined with a drawing which mount corresponds to which resonance we see in the spectrum. We will use this to start reducing (or even eliminating) those resonances starting with the most dominant ones close to 1kHz

Attached a copy of the drawing.



  818   Fri Feb 10 23:29:39 2012 frank, taraDailyProgressNoiseBudgetidentifying mechanical peaks in the beat noise

We try to identify the origins of each mechanical peaks in the beat signal from frequency between 100 up to a few kHz, so that we can damp it down below coating thermal noise. For today, we get rid of the peaks from the beam splitter in the beat path.


fig1: beat signal on linear x scale. Blue and red was taken with 10kHz and 5 kHz input range respectively. The broad band noise on Blue is from LO phase noise.


The beam splitter with a steel post and a clamp is replaced by ->


this block which holds the beam splitter inside. The whole pattern does not line up with the table, so we have to use clamps for now.


  817   Thu Feb 9 22:32:51 2012 FrankNotesBEATcollection of some numbers for future calculations

mixers  (which we currently have (and use) in the lab):

  1. ZFM-3-S (datasheet)
    - 0.04 to 400 MHz
    - Level 7
    - conversion loss @160MHz: 4.79dB typ
    - VSWR RF Port : 1.15:1
    - VSWR LO Port : 2.46:1
    - 1 dB COMP.: +1 dBm typ.
  2. ZX05-1MHW (datasheet)
    - 0.5 to 600 MHz
    - Level 13
    - conversion loss @160MHz: 5.2dB typ
    - VSWR RF Port : 1.17:1
    - VSWR LO Port : 2.37:1
    - 1 dB COMP.: +9 dBm typ.

cables  calculator for cable loss, which has a huge amount of different cable types in it's database. : http://vk1od.net/calc/tl/tllc.php

velocity ~0.66 to 7*c

RG-58C/U: loss 32 dB for 500ft @160MHz
                    Delay  5ns/m

RG-142    loss 25 dB for 500ft @160MHz
                   Delay  4.7ns/m


4-way splitter ZBSC-413 (datasheet)

additional insertion loss 0.5dB
power Input: 1W max


ZHL-1A.pdf (datasheet)
 - gain 16dB
 - output power: 28dBm min. (1dB compression)
 - noise figure: 11dB
- VSWR 2:2 (in and out)
 - 2 to 500 MHz

ZFL-500LN (datasheet)
 - gain 24dB 
 - output power: 5dBm min. (1dB compression)
 - noise figure: 2.9dB
- VSWR 1.5:1 (in) and 1.8:1 (out)
 - 0.1 to 500 MHz


  816   Thu Feb 9 22:31:03 2012 frank, taraDailyProgressBEATnoise budget and beat

Noise calculation from PD in PLL: ( I actually asked Koji once and did this already, see psl:730 . The results are similar)

1) determining which setup gives the best performance:

  • Gain on SR560 = 200. This gives UGF of 33kHz with 60 degree phase margin. Gain 500 has phase margin of 18 degree, which is too low, see fig 1.
  • Tuning range on Marconi = 1kHz. Currently, we cannot go with lower range. Usually, this noise couple directly to the readout and cannot be suppressed, so the lower noise (smaller range) the better.


fig1: OLG TF of PLL with different gain setup.



2) Measure electronic noise from readout system with the chosen setup. This noise will show up (after some correction) in the beat and determine what is the limitation of PLL readout technique.

     The PD was blocked, the feedback signal (Vfb) to the actuator (LO) was removed and measured.

3) Block diagram

[add block diagram and calculation]

4) After Koji explained on how to calculated noise budget from electronic noise in PLL to us, here the nb with PLL noise. (note: the LO phase noise has updated to 1kHz input range)


With the electronic noise from PLL, the sensitivity of this technique will prevent us from observing coating noise above 1kHz.

I'll calculate the noise from cable delay technique later and compare which one will give us better sensitivity.


Attachment 2: nb_2012_02_09.png
  815   Wed Feb 8 22:10:56 2012 frank, taraDailyProgressBEATnoise budget and beat

The noise budget is updated and plotted with today's beat measurement.

(*The electronic noise plotted in the graph is not correctly calibrated, see psl:816 for the complete calculation)

     After we replaced the table's broken leg, we floated the table and measured the beat signal as a reference before modifying the seismic stack. The calculation agree with the measured data quite well.

   We also measured the electronic noise from PLL. This was the signal which was fed back to the LO (with SR560 gain = 20). Apparently, we are sitting on it at 1kHz and above. We definitely need to work on PLL readout system to measure at lower sensitivity.


New traces in the noise budget:

  1. Beat noise (floated table)
  2. PLL electronic noise

I removed noises from room temperature fluctuation/ heater noise and spacer thermal noise because they are way smaller than coating noise and crowd the plot.

==Details about some traces in the noise budget==

Seismic noise:

     The vertical seismic noise coupling is calculated by applying the seismic measurement times stack transfer function times cavity bending coupling

     [ Frequency noise from seismic ] = [measured data] x [stacks TF] x [bending factor].

  • For seismic noise, The data was taken by a seismometer on a floated table.
  • Stack TF is computed based on springs and the weight of the blocks, see PSL: .. 
  • The cavity bending coupling is obtained from COMSOL simulation. However, the value used in the noise budget is about 1 order of magnitude larger than the value COMSOL provides. This is probably due to the fact that the cavity's support positions do not match exactly at the optimum point.

The peak around 6 Hz might be coupled from horizontal direction. This will be added soon.

RIN induced length noise is still an estimated. We have not been able to measured the real coupling yet, as the SNR is so low.

LO phase noise:  This is from measurement. I'm not quite sure if I miss some calibration factors. The phase noise does not show up in the beat yet even though it is very close together right now.

  814   Tue Feb 7 21:55:05 2012 Tara, FrankDailyProgressSeismicRTV springs

We measured Q and k of RTV springs. Currently, there are 9 of them. The results are ok

An aluminum block with accelerometer (0.398kg) was placed on a spring. The block was tapped, and the ringdown response was measured.


piece freq [Hz] spring constant k (N/m) Q
1 38.1 22,808 13.2
2 35 19,248 13.1
3 35.2 19,468 11
4 38.5 23,290 14.5
5 36.5 20,933 13.8
6 37.0 21,510 13.9
7 37.2 21,743 13.5
8* 32 16,090 9.55
9** 31.5 15,591 9.8

* the piece is broken, so it is softer and more lossy.

** the piece is made of other material.


Mean Q (piece 1 to 7) = 13.29 +/- 1.11

Mean k (piece 1 to 7) = 21,286 +/- 1,500

  813   Tue Feb 7 20:10:34 2012 Tara, FrankDailyProgressScheduleupdate

list of finished items

  • in-vacuum cables and connectors, screws, washers, teflon pieces, sensors etc cleaned - baking over night
  • cut & cleaned RTV springs - baking over night
  • polished and cleaned second shield
  • measured spring constant for all RTV springs to compare later with stack TF - see next elog entry for details
  • got lots of parts from the machine shop (periscope, pd mount, parts for beat setup - all parts cleaned and ready for assembly

all parts needed for upgrading the stack and adding radiative shields/heaters are tested/ready or currently beeing baked. Installation can start Wed/Thu

unfinished things left for tomorrow:

  • replace leg
  • replace opamp in PDHbox

P1820549.JPG P1820550.JPG

  812   Mon Feb 6 23:52:24 2012 FrankNotesScheduleProject schedule for this week


  • clean and bake in-vacuum cables and connectors
  • clean other components, e.g screws
  • cut, clean and bake RTV springs
  • modify and fix ACAV servo
  • replace broken leg and measure beat


  • open chamber
  • measure TF from table to top of stack
  • replace springs
  • re-measure TF with new setup
  • add heaters/shield
  • close chamber and start pumping
  • build beat-breadboard (if all mounts are ready)


  • re-align cavities
  • replace input periscope
  • re-align cavities with new periscope
  • replace PD mounts
  • detailed heater characterization
  • try to measure TF from heater to cavity length


  • re-align everything
  • shift beat frequency down to 10MHz (?)
  • try feedback control to heater
  • measure detailed TF from heater to cavity length over weekend
  811   Mon Feb 6 22:35:49 2012 frank, taraDailyProgressElectronics Equipmentacav characterization

TF from all components except the sevo is measured (collectively). This will help us to determine what kind of servo(UPDH) we want for ACAV.


TF from other components in ACAV loop except the servo (collectively) (frequency discriminator, marconi, amplifier) and fit]


TF from ACAV loop without servo. The red curve is fitted with 1pole at 50kHz with 70kHz delay (exp(-i*f/70kHz)) . Since our bandwidth of interest extends upto only a few hundred kHz, this fit is good enough for a model. It starts to diverge from the data at 300kHz.

Note: The TF looks ok, it is flat as expected from most of the parts (frequency disc/Marconi/amplifier). The pole and time delay is from the AOM. We can see the phase changes as we change the AOM position so that the beam is closer to the PZT side. We gain ~5 degrees from adjusting the position.

[current updh schematic]

The TF has a pole around 50kHz. C18 with 3300pF gives a zero at that frequency and cancels the pole. We are designing the TF of the servo that is suitable for our need (UGF ~100kHz, with 1/f roll off at UGF, and ~45 degree phase margin, and, 1/f^2 at low frequency)

Problem with UPDH, on stage 2, the resistors' values might be wrong. We will check and fix it tomorrow.



  810   Sat Feb 4 03:18:08 2012 taraDailyProgressElectronics Equipmentacav characterization

I checked the TF of the amplifier used in the ACAV loop because I did not measure that yesterday.  The amplifier's TF is flat at least up to 200 kHz. The Bode plots between the loop with and without the amplifier are pretty much the same. Thus, it is ok and won't cause a problem in the loop.


The setup is similar to what I did in the previous entry, except the output of the marconi is connected with the amplifier. See blue arrow.





The TFs between the two cases have similar shape. So the amplifier will not be the limiting component. The magnitudes (in arbitrary unit) are slightly different because I did not attenuate the power by the same factor it was amplified.


The UPDH box will be modified next to see if we can increase the loop bandwidth.

  809   Thu Feb 2 23:51:32 2012 frank, taraDailyProgressElectronics Equipmentacav characterization

We measured the TF of marconi using PLL loop. Marconi has flat response up to around 200kHz. This is quite good and we can certainly use it in ACAV loop.

 ==block diagram of ACAV loop==

  • D: frequency disc
  • G: servo
  • A: Marconi
  • H: Amplifier


The whole OLG TF was measured in PSL:...  This time we looked into the marconi to see if its TF has bandwidth high enough for ACAV loop or not.  We know that Marconi has lower phase noise than LIGO homemade VCO (Megan's elog), but we have not learned about its bandwidth yet.

==PLL setup for Marconi TF==

The actual magnitude TF at DC can be determined by using a voltage calibrator to inject DC signal and measure how much the frequency of the output changes. This depends on the tuning range setup on the Marconi. However, we don't know the bandwidth of the TF, so we use PLL to find out. The setup is shown below.  The gain from SR560 was set to be low, so that the signal at high frequency will be the TF of the Marconi + frequency discriminator.



Since the mixer output gives 1/f response (flat in [rad/rtHz] unit), we corrected the TF by multiplying back with f to get the TF of the marconi. The magnitude on the plot below has arbitrary unit. We are only interested in its shape. We tried 100Hz, 1kHz(not shown), 10kHz tuning range. The magnitude varies with the tuning range as expected. The phase does not change that much.



If we want to have phase margin of 45 degree, assuming other components in the loop have no phase lag. The best UGF we can do is upto 200kHz, according to the phase response of the Marconi (the phase drops by 135 degree around 200 kHz). Therefore, using a Marconi as an oscillator for driving the AOM is also possible because its bandwidth is high enough for measurement up to 1kHz.

Note: we will check what is the TF of the amplifier (H) used in our setup to make sure that it is not the limiting component.

We can definitely use Marconi as a VCO in our ACAV loop.

NExt: The next step is checking the UPDH box. At a glance, we found that the TF shape of the current UPDH is not suitable for our requirement.

  808   Wed Feb 1 15:46:51 2012 taraDailyProgressElectronics Equipmentacav characterization

I'm working on characterizing ACAV loop to check if the loop is good enough for the coating noise measurement. The results show that there we need better improvement on this part.


I realigned the cavity, and the visibility is only 85%. I have not figured out yet why the visibility decreases from 90%. Then I measured dark noise @ error point, inloop noise with different gain setting, and error signal's slope ( all with 50 ohm system).


 Error slope (with 2V DC on RFPD) =122.7 kHz. The visibility is @ 85%

1) Dark noise and inloop noise:  The in loop noise changes with gain setup on UPDH box. I changed the gain from 3 to 7 and made sure that there was no oscillation (oscillating @ gain7.5).


 With the error signal slope, I can calibrate in loop noise to absolute frequency noise from ACAV loop and plot them on noise budget. On the figure below, we can see that at high frequency noise from ACAV loop is dominating.


Notice the bump around 3kHz. It comes from the phase noise of the marconi. If I change the tuning range from 10khz to 1kHz the in loop noise also change (blue to green in the below figure). It is a trade off between lower frequency noise, less gain.   The peaks around 80 Hz go up because we have less gain to suppress the noise.



2) OLG TF at different gain setup



Back to the 10kHz input range. Since the in loop noise is getting worse with more gain, I look into the OLG TF of the loop. The UGF is around 3-10kHz, depending on the gain. TF looks ok at gain 3-5, the magnitude increases along with the gain. At gain 7, the TF starts to deviate at high frequency (it might be oscillating at high frequency).

The in loop noise should be suppressed with 1/(1+G) factor, but the results above do not tell the same story. The gain increases but the in loop noise gets worse. So the next step is to look deeper into this problem. I'll use SR650 instead of UPDH and check if the same problem occur or not. This should verify if the UPDH box is bad or not.

Note: I'll calculate what we need for ACAV loop, and check if LIGO homemade VCO will be good enough for ACAV loop or not.

  807   Tue Jan 31 01:26:03 2012 taraNotesDrawingslegs and base for RFPD

These drawings are for RFPD's bases. The square plate will be made of Aluminum. It is for mounting RFPD on the table. The block is for RFPD legs which will be made from Delrin.

Attachment 1: RFPD_base.PDF
Attachment 2: RFPD_leg.PDF
  806   Mon Jan 30 22:53:30 2012 frank, taraDailyProgressNoiseBudgetLoop characterization

swap the RFPDs. The SN002 for RCAV was not working (see psl:803). So we use SN01 for RCAV and characterize it again.

1) dark noise in Vrms/  rt Hz


Dark noise is not particularly high. ~40 nV/rt Hz flat. In loop noise is only around 10 nV/rt Hz. It is about a factor of 2 above SR785 noise for the input range setup. When I measured the dark noise, I made sure that monitor screens in the lab were off, no cables connected to a scope to create any ground loop. So the measured signal this time does not have harmonic lines as much as the previous measurement does.

2) error signal slope :  measured to be 34.4 kHz/V ( with 20dB attenuator on EOM. if I use 16dB attenuator, the calibration can be lower to 22 kHz/V). We will decide what should be the modulation index once we calculate the acceptable level of shot noise and compare with the gain we need.

Note that the numbers in psl:802 missed a factor of 2 (they should be twice as much).

3) absolute frequency noise from dark noise at error point and inloop noise: The gain setup for common/fast are 1000/750. It becomes unstable if I increase more fast gain. 


 The inloop noise lies below the estimated coating noise up to 100kHz. So if we can decrease any other technical noise sources, we should be able to reach coating noise level.


Next will be the complete OLG TF measurement.

  805   Sat Jan 28 02:41:45 2012 FrankPhotosopticphotos of cavity input optics

just for reference which part is/was where for later...

IMG009.jpg IMG010.jpg

IMG011.jpg IMG012.jpg

  804   Sat Jan 28 02:24:07 2012 FrankNotesDrawingsPeriscope checked

Double checked the distance between the mirrors using the new periscope base Tara designed. Distance is 2.96in according to the 3D assembly, close enough to the 3" we want. CAD files are on the svn in the mechanical drawings folder.


  803   Fri Jan 27 22:24:59 2012 FrankNotesElectronics EquipmentRCAV RFPD broken

Took me some time to figure out that the current RCAV RFPD  (SN002) is broken, but not in a normal way. The functionality is still as it should be using standard measurements: RF-noise level is low, response is OK etc. , but some non-stationary noise process at RF frequencies happens from time to time, only visible in the demodulated spectrum. My guess is that the MAX4107 is about to die or close to be unstable and starts oscillating or so. Will check. ACAV PD (SN001) is fine.

Attached a movie of the demodulated noise spectrum. The "feature" does not show up on the oscilloscope. There you see only constant noise, no changes in amplitude or so. Only in the demodulated spectrum shows the effect, mostly visible between 10k and 100k. Rana, Koji: have you guys seen this before?


  802   Thu Jan 26 21:05:21 2012 frank, taraDailyProgressNoiseBudgetLoop characterization

We need to characterize RCAV and ACAV loop, here is a list to do

  1. Power adjustment
  2. RCAV loop characterization
  3. ACAV loop characterization

Power adjustment: We now choose some appropriate and easier numbers of the setup for future reference.

The DC out from RCAV and ACAV RFPD are 2 V which correspond to ~ 1mW input.

The DC out from RFAM PD is 0.2 V which corresponds to 1mW input as well (this one has different R, hence different gain)

[add pic], and power to EOM is 0dBm. This will be our current standard setup.

RCAV loop characterization:

We use different RFPD, different TTFSS set, so this has to be done again.

The power to EOM is 0 dBm, 1mW input to RCAV:

The error signal slope:

ACAV 0.0149 MHz/V (I'm not sure why the error signal from ACAV is clipped,see below picture)

RCAV 0.0224 MHz/V

[add fig]

Dark noise level in Vrms/ rtHz


-> Dark noise level in absolute frequency:


 It looks like the noise is low enough almost up to 1kHz. Yay!

The OLG transfer function as well as in loop noise, will be measured later once we optimize the gain.

ELOG V3.1.3-